# Exploring the Vertical-Domain Reasoning Capabilities of Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) are reshaping learning paradigms, cognitive processes, and research methodologies across a wide range of domains. Integrating LLMs with professional fields and redefining the relationship between LLMs and domain-specific applications has become a critical challenge for promoting enterprise digital transformation and broader social development. To effectively integrate LLMs into the accounting domain, it is essential to understand their domain-specific reasoning capabilities. This study introduces the concept of vertical-domain accounting reasoning and establishes evaluation criteria by analyzing the training data characteristics of representative GLM-series models. These criteria provide a foundation for subsequent research on reasoning paradigms and offer benchmarks for improving accounting reasoning performance. Based on this framework, we evaluate several representative models, including GLM-6B, GLM-130B, GLM-4, and OpenAI GPT-4, on a set of accounting reasoning tasks. Experimental results show that different prompt engineering strategies lead to varying degrees of performance improvement across models, with GPT-4 achieving the strongest accounting reasoning capability. However, current LLMs still fall short of real-world application requirements. In particular, further optimization is needed for deployment in enterprise-level accounting scenarios to fully realize the potential value of LLMs in this domain.
Exploring the Vertical-Domain Reasoning Capabilities of Large Language Models
@@ 1 Introduction
The main contributions of this paper are as follows. First, we extend research on general-purpose LLMs to vertical-domain accounting reasoning. We analyze the relationships between LLM reasoning ability, mathematical reasoning, logical inference, and accounting expertise, providing a theoretical foundation for subsequent studies. Second, we design a comprehensive and diverse benchmark dataset for accounting reasoning evaluation. Unlike existing datasets that focus on single financial subdomains, our dataset covers accounting principles, financial statements, cost accounting, and auditing, enabling a more comprehensive assessment of LLM accounting reasoning performance. Third, we evaluate representative LLMs, including GPT-4 and GLM-series models, on the proposed benchmark and analyze their strengths and limitations, offering valuable insights for accounting-related applications. Finally, we explore the integration of domain knowledge with reasoning optimization strategies to improve LLM performance in accounting tasks, providing guidance for future research and accelerating the deployment of intelligent accounting systems in real-world practice.
@@ 2.1 Development of LLMs in the Accounting Domain
The successful application of general-purpose large language models (LLMs) in vertical domains such as medicine and law provides important references for their deployment in the accounting domain. In medical applications, LLMs have been used in clinical decision support, medical image interpretation, and medical record generation (panagiotou2024). Due to the high interpretability of medical imaging data, LLMs can effectively identify visual patterns and, when combined with natural language generation, improve the transparency and explainability of AI systems, leading to better clinical outcomes (chalkidis2023).
@@
Existing research mainly focuses on the following directions. First, some studies aim to improve specific reasoning abilities of LLMs, such as mathematical or logical reasoning (huang2022reasoning). However, these methods often struggle to integrate accounting-specific knowledge, as they assume that reasoning capability can be enhanced independently of domain expertise. Consequently, such approaches show limited effectiveness in accounting scenarios (li2023). Moreover, there remains a lack of work explicitly targeting accounting reasoning.
Second, some studies attempt to improve domain performance by specializing training data or scaling model parameters. Models such as BloombergGPT (wu2023bloomberggpt) and FinGPT (liu2023fingpt) incorporate large-scale financial corpora or increase model capacity to enhance performance. However, these approaches do not necessarily lead to stable improvements in reasoning ability and often lack interpretability, making it difficult to assess whether performance gains stem from genuine reasoning improvements or surface-level pattern matching (zhao2023).
Third, evaluation in some works primarily focuses on answer accuracy. For example, LLMs are evaluated on tasks such as financial question answering or accounting knowledge retrieval (faturos2023; theuma2024; shah2023). These tasks are largely text-based and do not reflect the structured, multi-step nature of accounting reasoning. As a result, they fail to capture the core reasoning capabilities required in accounting applications.
Fourth, several exploratory studies discuss the potential of LLMs in accounting practice (he2023; ouyang2024). While these works provide valuable insights, they have not yet established systematic research paradigms or concrete optimization strategies for improving accounting reasoning capabilities.
@@ 2.2 Reasoning Capability of LLMs in Accounting
In LLM research, reasoning is commonly defined as the ability to infer conclusions from given facts and constraints (huang2022reasoning; yu2023). Logical reasoning evaluations typically assess the correctness of model outputs on inference tasks. According to reasoning structure, reasoning can be categorized into mathematical reasoning, logical reasoning, causal reasoning, and commonsense reasoning (sun2023; huang2022reasoning). Among these categories, mathematical reasoning and logical reasoning are most closely related to accounting tasks.
Accounting problems often involve complex numerical computation, multi-step logical deduction, and strict rule-based constraints. Therefore, reasoning capability plays a central role in determining LLM performance in accounting applications. Logical reasoning evaluates whether models can derive valid conclusions based on a set of premises and rules, typically measured through natural language inference tasks (srivastava2022; yu2023; chang2023). However, existing inference benchmarks are designed for general language understanding and do not adequately reflect the structured reasoning demands of accounting.
Mathematical reasoning evaluates a model’s ability to perform arithmetic and symbolic computation. This capability is commonly assessed using datasets such as GSM8K (cobbe2021), which consist of multi-step elementary math word problems. These tasks require sequential reasoning and intermediate result propagation, which closely resemble accounting calculation processes.
In accounting contexts, except for high-level financial analysis, most tasks involve multi-step numerical computation. If LLMs fail to correctly propagate intermediate results across reasoning steps, errors accumulate and lead to incorrect conclusions. Therefore, multi-step mathematical reasoning is particularly critical. Moreover, many accounting scenarios require reasoning over temporally evolving entities and constraints, where integrating temporal knowledge graphs (xiongtilp; xiong2024teilp; xiong2024large) with LLM-based symbolic reasoning frameworks (yang2024harnessing) can provide structured state tracking and improve consistency across multi-step inference. SWAP (xiong2025deliberate) introduces structure-aware planning to enhance multi-step reasoning performance. MR-GSM8K (zeng2024) extends GSM8K by filtering problems with short reasoning chains and increasing reasoning depth, resulting in longer and more complex reasoning trajectories, resulting in longer and more complex reasoning trajectories that better resemble real-world accounting calculations. After further filtering, 586 MR-GSM8K problems requiring at least three computation steps remain. This subset can serve as a foundational benchmark for evaluating multi-step calculation ability in accounting contexts, referred to as the Multi-Calculation Benchmark.
@@ 2.3 Evaluation Benchmarks for Chinese Accounting Reasoning
The Chinese Language Understanding Evaluation (CLUE) benchmark (xu2020) is a widely used Chinese counterpart to GLUE (wang2018). CLUE includes multiple natural language understanding tasks such as text classification, reading comprehension, and natural language inference, enabling comprehensive evaluation of Chinese language understanding. However, CLUE primarily targets general language understanding and does not explicitly address accounting reasoning.
@@ 3.1 Experimental Setup
Empirical studies from the Stanford Alpaca team indicate that using curated datasets at approximately 5 percent of the original size can significantly improve model performance (xia2024). High-quality data plays a critical role in enhancing both reasoning accuracy and robustness. Accordingly, the first step of our evaluation is to assess the baseline reasoning capabilities of LLMs using general-purpose reasoning benchmarks, followed by specialized evaluation on accounting-oriented datasets.
@@ 3.2 Evaluation of Mathematical Reasoning Ability
Existing benchmarks are widely used to evaluate the mathematical reasoning abilities of LLMs, including natural language inference and multi-step arithmetic reasoning datasets (srivastava2022; yu2023; chang2023). However, most general-purpose benchmarks are not designed to capture the complexity of accounting computation, particularly under real-world constraints such as conditional logic and cumulative error propagation.
Among existing datasets, GSM8K (cobbe2021) is one of the most widely used benchmarks for evaluating multi-step mathematical reasoning. It consists of 8.5K high-quality elementary math word problems that require two to eight reasoning steps, primarily involving arithmetic operations such as addition, subtraction, multiplication, and division. These problems resemble simplified accounting calculations, where each intermediate result must be correctly derived to obtain the final answer.
However, GSM8K contains a large number of short problems with limited reasoning depth. To better evaluate extended reasoning chains, MR-GSM8K (zeng2024) is constructed by filtering GSM8K problems with short reasoning trajectories and retaining those that require longer multi-step computation. Problems with answers shorter than 300 words are removed, resulting in a dataset that emphasizes sustained reasoning and error accumulation.
@@ 3.3 Evaluation of Chinese Accounting Reasoning Ability
The Chinese Language Understanding Evaluation (CLUE) benchmark (xu2020) is widely used to evaluate general Chinese language understanding and reasoning capabilities. It includes tasks such as text classification, reading comprehension, and natural language inference, and serves as the Chinese counterpart to GLUE (wang2018). However, CLUE primarily evaluates general reasoning rather than accounting-specific inference.

### v2

# Accounting Reasoning in Large Language Models: Concepts, Evaluation, and Empirical Analysis

@@ Abstract
Large language models (LLMs) are increasingly reshaping learning paradigms, cognitive processes, and research methodologies across diverse domains. As their adoption expands, effectively integrating LLMs into professional fields and clarifying their role in domain-specific applications has become a key challenge for enterprise digital transformation and broader societal development. In the accounting domain, successful integration requires a systematic understanding of LLMs’ domain-specific reasoning capabilities. In this study, we introduce the concept of accounting reasoning and propose a set of evaluation criteria grounded in an analysis of the training data characteristics of representative GLM-series models. These criteria establish a foundation for studying accounting-oriented reasoning paradigms and provide benchmarks for assessing and improving model performance. Building on this framework, we evaluate several representative LLMs, including GLM-6B, GLM-130B, GLM-4, and GPT-4, across a range of accounting reasoning tasks. Our experimental results show that prompt engineering strategies can yield varying degrees of performance improvement across models, with GPT-4 demonstrating the strongest overall accounting reasoning capability. Nevertheless, the results indicate that current LLMs remain insufficient for real-world accounting applications. In particular, further optimization is required for deployment in enterprise-level accounting scenarios to fully realize the potential value of LLMs in this domain.
Accounting Reasoning in Large Language Models: Concepts, Evaluation, and Empirical Analysis
@@ 1 Introduction
The main contributions of this paper are as follows. First, we extend research on general-purpose LLMs to accounting reasoning. We analyze the relationships between LLM reasoning ability, mathematical reasoning, logical inference, and accounting expertise, providing a theoretical foundation for subsequent studies. Second, we design a comprehensive and diverse benchmark dataset for accounting reasoning evaluation. Unlike existing datasets that focus on single financial subdomains, our dataset covers accounting principles, financial statements, cost accounting, and auditing, enabling a more comprehensive assessment of LLM accounting reasoning performance. Third, we evaluate representative LLMs, including GPT-4 and GLM-series models, on the proposed benchmark and analyze their strengths and limitations, offering valuable insights for accounting-related applications. Finally, we explore the integration of domain knowledge with reasoning optimization strategies to improve LLM performance in accounting tasks, providing guidance for future research and accelerating the deployment of intelligent accounting systems in real-world practice.
@@ 2.1 Development of LLMs in the Accounting Domain
The successful application of general-purpose large language models (LLMs) in vertical domains such as medicine and law provides important references for their deployment in the accounting domain. In medical applications, LLMs have been used in clinical decision support, medical image interpretation, and medical record generation (Yang et al., 2023). Due to the high interpretability of medical imaging data, LLMs can effectively identify visual patterns and, when combined with natural language generation, improve the transparency and explainability of AI systems, leading to better clinical outcomes (Guha et al., 2023).
@@
Existing research mainly focuses on the following directions. First, some studies aim to improve specific reasoning abilities of LLMs, such as mathematical or logical reasoning (Huang and Chang, 2023). However, these methods often struggle to integrate accounting-specific knowledge, as they assume that reasoning capability can be enhanced independently of domain expertise. Consequently, such approaches show limited effectiveness in accounting scenarios (Gu et al., 2025). Moreover, there remains a lack of work explicitly targeting accounting reasoning.
Second, some studies attempt to improve domain performance by specializing training data or scaling model parameters. Models such as BloombergGPT (Wu et al., 2023) and FinGPT (Wang et al., 2023) incorporate large-scale financial corpora or increase model capacity to enhance performance. However, these approaches do not necessarily lead to stable improvements in reasoning ability and often lack interpretability, making it difficult to assess whether performance gains stem from genuine reasoning improvements or surface-level pattern matching (Diaz and Madaio, 2024).
Third, evaluation in some works primarily focuses on answer accuracy. For example, LLMs are evaluated on tasks such as financial question answering or accounting knowledge retrieval (Srivastava et al., 2024; Xie et al., 2024; Zhang et al., 2023). These tasks are largely text-based and do not reflect the structured, multi-step nature of accounting reasoning. As a result, they fail to capture the core reasoning capabilities required in accounting applications.
Fourth, several exploratory studies discuss the potential of LLMs in accounting practice (Vasarhelyi et al., 2023; Zhao and Wang, 2024). While these works provide valuable insights, they have not yet established systematic research paradigms or concrete optimization strategies for improving accounting reasoning capabilities.
@@ 2.2 Reasoning Capability of LLMs in Accounting
In LLM research, reasoning is commonly defined as the ability to infer conclusions from given facts and constraints (Huang and Chang, 2023; Mondorf and Plank, 2024). Logical reasoning evaluations typically assess the correctness of model outputs on inference tasks. According to reasoning structure, reasoning can be categorized into mathematical reasoning, logical reasoning, causal reasoning, and commonsense reasoning (Plaat et al., 2024; Huang and Chang, 2023). Among these categories, mathematical reasoning and logical reasoning are most closely related to accounting tasks.
Accounting problems often involve complex numerical computation, multi-step logical deduction, and strict rule-based constraints. Therefore, reasoning capability plays a central role in determining LLM performance in accounting applications. Logical reasoning evaluates whether models can derive valid conclusions based on a set of premises and rules, typically measured through natural language inference tasks (Srivastava et al., 2023; Mondorf and Plank, 2024; Fan et al., 2024). However, existing inference benchmarks are designed for general language understanding and do not adequately reflect the structured reasoning demands of accounting.
Mathematical reasoning evaluates a model’s ability to perform arithmetic and symbolic computation. This capability is commonly assessed using datasets such as GSM8K (Cobbe et al., 2021), which consist of multi-step elementary math word problems. These tasks require sequential reasoning and intermediate result propagation, which closely resemble accounting calculation processes.
In accounting contexts, except for high-level financial analysis, most tasks involve multi-step numerical computation. If LLMs fail to correctly propagate intermediate results across reasoning steps, errors accumulate and lead to incorrect conclusions. Therefore, multi-step mathematical reasoning is particularly critical. Moreover, many accounting scenarios require reasoning over temporally evolving entities and constraints, where integrating temporal knowledge graphs (Xiong et al., ; Xiong et al., 2024b, a) with LLM-based symbolic reasoning frameworks (Yang et al., 2024) can provide structured state tracking and improve consistency across multi-step inference. SWAP (Xiong et al., 2025) introduces structure-aware planning to enhance multi-step reasoning performance. MR-GSM8K (Zeng et al., 2023) extends GSM8K by filtering problems with short reasoning chains and increasing reasoning depth, resulting in longer and more complex reasoning trajectories, resulting in longer and more complex reasoning trajectories that better resemble real-world accounting calculations. After further filtering, 586 MR-GSM8K problems requiring at least three computation steps remain. This subset can serve as a foundational benchmark for evaluating multi-step calculation ability in accounting contexts, referred to as the Multi-Calculation Benchmark.
@@ 2.3 Evaluation Benchmarks for Chinese Accounting Reasoning
The Chinese Language Understanding Evaluation (CLUE) benchmark (Xu et al., 2020) is a widely used Chinese counterpart to GLUE (Wang et al., 2018). CLUE includes multiple natural language understanding tasks such as text classification, reading comprehension, and natural language inference, enabling comprehensive evaluation of Chinese language understanding. However, CLUE primarily targets general language understanding and does not explicitly address accounting reasoning.
@@ 3.1 Experimental Setup
Empirical studies from the Stanford Alpaca team indicate that using curated datasets at approximately 5 percent of the original size can significantly improve model performance (Li et al., 2024). High-quality data plays a critical role in enhancing both reasoning accuracy and robustness. Accordingly, the first step of our evaluation is to assess the baseline reasoning capabilities of LLMs using general-purpose reasoning benchmarks, followed by specialized evaluation on accounting-oriented datasets.
@@ 3.2 Evaluation of Mathematical Reasoning Ability
Existing benchmarks are widely used to evaluate the mathematical reasoning abilities of LLMs, including natural language inference and multi-step arithmetic reasoning datasets (Srivastava et al., 2023; Mondorf and Plank, 2024; Fan et al., 2024). However, most general-purpose benchmarks are not designed to capture the complexity of accounting computation, particularly under real-world constraints such as conditional logic and cumulative error propagation.
Among existing datasets, GSM8K (Cobbe et al., 2021) is one of the most widely used benchmarks for evaluating multi-step mathematical reasoning. It consists of 8.5K high-quality elementary math word problems that require two to eight reasoning steps, primarily involving arithmetic operations such as addition, subtraction, multiplication, and division. These problems resemble simplified accounting calculations, where each intermediate result must be correctly derived to obtain the final answer.
However, GSM8K contains a large number of short problems with limited reasoning depth. To better evaluate extended reasoning chains, MR-GSM8K (Zeng et al., 2023) is constructed by filtering GSM8K problems with short reasoning trajectories and retaining those that require longer multi-step computation. Problems with answers shorter than 300 words are removed, resulting in a dataset that emphasizes sustained reasoning and error accumulation.
@@ 3.3 Evaluation of Chinese Accounting Reasoning Ability
The Chinese Language Understanding Evaluation (CLUE) benchmark (Xu et al., 2020) is widely used to evaluate general Chinese language understanding and reasoning capabilities. It includes tasks such as text classification, reading comprehension, and natural language inference, and serves as the Chinese counterpart to GLUE (Wang et al., 2018). However, CLUE primarily evaluates general reasoning rather than accounting-specific inference.